OpenStudy (anonymous):

Why is the lim x->0+ of (1/x)/(-1/x^2) = lim x->0+ (-x) = 0 Makes no sense to me! Why is the lim x->0+ of (1/x)/(-1/x^2) = lim x->0+ (-x) = 0 Makes no sense to me! @Mathematics

OpenStudy (anonymous):

limx->0+ of (1/x) is infinity limx->0+ of (-1/x^2) is -infinity So WHY does infinity/-infinity = (-x) or 0 now....

OpenStudy (anonymous):

Makes no goddam sense...

OpenStudy (anonymous):

If:\[x\neq 0\]then: \[\frac{\frac{1}{x}}{-\frac{1}{x^2}}\cdot \frac{x^2}{x^2}=\frac{x}{-1}=-x\]

OpenStudy (anonymous):

what... how am i supposed to know to do that!!!

OpenStudy (anonymous):

you have this big fraction with small fractions in the numerator and denominator. Any time you have that, you almost always want to get rid of the internal fractions.

OpenStudy (anonymous):

Okay, will keep that in mind! Thanks for the help!

OpenStudy (anonymous):

Also note, its only correct to think about dividing limits if the limits actually exist. So because both:\[\lim_{x\rightarrow0}\frac{1}{x}\]and \[\lim_{x\rightarrow 0}\frac{1}{x^2}\]dont exist, try not to think about division.

OpenStudy (anonymous):

Well i'm doing l'hopitals, so I have to think about division, right?. That was just one step from original function xlnx in my book that I did not understand. They didn't mention at all that they multiplied by x^2 so I was very confused.

OpenStudy (anonymous):

ah, i see. yeah, youre right.

OpenStudy (anonymous):

Joe, I have one more quick question. I feel they are going to trick me tomorrow on the quiz giving me a question that looks like it needs L'Hoptals, but actually doesn't. The first conditions for lhopitals is that f and g be differentiable. Which implies continuity. Which I understand. But then my book says that \[g'(x)\neq0\] near a. Shouldnt it be \[f'(x)\neq0 AND g'(x)\neq0\] Why did they only mention g'(x)?

OpenStudy (anonymous):

I know all the intermediate form types, so don't worry about that. It's just that one condition I'm a bit unclear on.